. The demand for a monopolists product is given by Q monopolists marginal cost is given by MC -3 The profis-ma quantity of quantity output for this monopolist is A) 10o C) 40 8) 4444 D) a0 There is a payoff matrix of two flems: their collusion or competition (answer 14-15) differens profits are listed when they choose firm B competition firm A competition! 14.14 27.5 collusion 5. a8 9.19 2. In the game above, who has dominant strategy A) Player A has a dominant strategy c) Both players have dominant strategies. D) Neither player has a dominant strategy 3. Which one is the Nash equilibrium? A) firm A competition, firm B competition. B) firm A competition, firm B collusion B) Player B has a dominant strategy C rirm A collusion, firm B competition.D) firm A collusion, firm B collusion. In the following game, all payoffs are listed with the row players payoffs first and the column players payoffs second (19-20) Player B BI B2 Player A A1 5,6 7,2 A2 4,5 9,1 14. In game above, a) Player A choosing Au and Player B choosing Bı is a Nash equilibrium. b) Player A choosing Az and Player B choosing B2 is a Nash equilibrium. e) there is no Nash equilibrium. d) there are multiple Nash equilibria in pure strategies. 5. In game above, a) Player A has a dominant strategy b) Player B has a dominant strategy c) Both players have dominant strategies d) Neither player has a dominant strategy Part 2: Short Questions (You need to show steps to get full credits) 1. Meimei Hans utility function is U-ioX Z, The price of X is sio, the price of Z is s
Answer #1

11. Profit is maximized where MR = MC

P = (400 - Q)/4

MR = d(PQ)/dQ = 100 - Q/2 = MC = 2Q

Q = 40

12. Both players have dominant strategies (both would choose to compete no matter what the other player chooses to do)

13. option A is correct

14. option A is correct (player A chooses A1 and player B chooses B1 - dominant strategy)

15. option B - player B has a dominant strategy (B earns a higher payoff using this strategy no matter what A chooses)